Q-A box of mass 9.6kg is at rest on level ground. What is the normal force acting on the block (in unit of N)?

Q2-If a block of mass 692g is being pulled with a force of 1.7N parallel to a horizontal surface with a max static friction coefficient of 0.999, but the block remains at rest what is the static friction force on the block (in unit of N)?

Q3-

A block of mass 2.25kg is being pulled by a string parallel to the contact surface with a slowly increasing force. If the block remains stationary until the force reaches 13.9N, what is the static friction coefficient?
Answer #1

q1) N-mg =0 as box is at rest a=0 F net = ma =0 so N-mg =0

N =mg = 9.6 *9.8 = 94.08 N